OpenStudy (anonymous):

9 1/2 to 21 3/8 write the ratio as a ratio of whole numbers using fractional notation

hero (hero):

\[\left( 9 \frac{1}{2}\right) \div \left( 21 \frac{3}{8}\right)\]